A cement block accidentally falls from rest from the ledge of a 53.4-m-high building. When the block is 14.2 m above the ground, a man, 1.91 m tall, looks up and notices that the block is directly above him. How much time, at most, does the man have to get out of the way?
Answer #1

DATA:

v_{0}=0

y=53.4\, m

y_{2}=14.2\, m

y_{m}=1.91\, m

g=9.81\, m/s^{2}

t=?

Solution:

see the following figure:

First, we calculate the distance y_{1} the block since released until the man sees it

y_{1}=y-y_{2}

y_{1}=53.4\, m-14.2\, m

y_{1}=39.2\, m

Now, we calculate the speed of the block v_{1} when it has fallen 39.2 m:

v_{1}^{2}=v_{0}^{2}+2.g.y_{1}

v_{1}=\sqrt{v_{0}^{2}+2.g.y_{1}}

Replacing values given in data:

v_{1}=\sqrt{(0)^{2}+2.(9.81\, m/s^{2}).(39.2\, m)}

v_{1}=27.7\, m

Second, we calculate the distance y_{3} down the block to the man's head:

y_{3}=y_{2}-y_{m}

y_{3}=14.2\, m-1.91\, m

y_{3}=12.3\, m

Now, we calculate the speed of the block v_{2} when it has fallen a distance y_{3}:

{v_{2}}^{2}={v_{1}}^{2}+2.g.y_{3}

v_{2}=\sqrt{{v_{1}}^{2}+2.g.y_{3}}

Replacing values given in data:

v_{2}=\sqrt{(27.7\, m/s)^{2}+2.(9.81\, m/s^{2}).(12.3\, m)}

v_{2}=31.8\, m/s

Finally, the time (at most) does the man have to get out of the way is calculated using:

v_{2}=v_{1}+g.t

Isolating t:

t=\frac{v_{2}-v_{1}}{g}

Replacing values:

t=\frac{31.8\, m/s-27.7\, m/s }{ 9.81\, m/s^{2}}

{\color{Blue} t=0.418\: s}